# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v2

@@ 4. Method
An overview of the training and densification process can be found in Figure 7, where we enumerate the individual steps of our training and densification process: To train models exceeding GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view. To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees have been replaced by SPTs, forming a hierarchical SPT (HSPT). To minimize costly transfers between RAM and VRAM, we keep track of which SPTs are currently loaded and at which detail, and cache them in GPU memory. Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM. Densification occurs on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes. The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations.
4.1. Initialization
Following Kerbl et al. (2024), we initialize the Gaussian model from a sparse point cloud, augmented with skybox points. This initial representation is small enough to fully reside in GPU memory and is trained for 100 000100000100\,000100 000 iterations with densification disabled. The goal of this phase is to establish a stable global scene structure before hierarchy construction. After this initial optimization, we build a binary Gaussian hierarchy, with the trained Gaussians as leaf nodes and parent nodes representing merged approximations of their children.
4.2. Gradient Propagation
To update the properties of Gaussians, including base color, SH coefficients, position, and covariance, we follow the standard 3DGS error propagation (Kerbl et al., 2023). However, we operate only on the cut set chosen for each training view. Thus, gradients may, in general, be propagated to Gaussians in the middle of the hierarchy.
4.3. Densification
@@
Densifying a Level-of-Detail representation presents a unique challenge: the underlying hierarchical structure must evolve continuously during training. Prior works circumvent this issue by constructing LoD hierarchies only after chunk-level training and densification are complete.
@@
We take inspiration from 3DGS-MCMC (Kheradmand et al., 2024), which selects Gaussians probabilistically (weighted by opacity) to be ‘split’, replacing it by two new Gaussians which together should appear similarly to the original split Gaussian. Noteably, this resembles the way a parent node in a Gaussian hierarchy approximates its children. Therefore, we adopt this approach and ‘spawn’ two new child nodes for a leaf instead of splitting a Gaussian, increasing the size of the hierarchy with minimal artifacts.
@@
Instead of pruning, Kheradmand et al. (2024) declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian. We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified. See Figure 3 for an overview of the two hierarchy densification operations. Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required. Avoiding destructive resets, as advocated by Kerbl et al. (2023), is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
Figure 3. Example of densifying and respawning leaf nodes.
4.4. The hierarchical SPT datastructure
@@
To enable GPU-accelerated cuts, Kerbl et al. (2024) evaluate the cut condition in parallel for each Gaussian:
$\Big{(}m_{d}(i)<\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big{)}\land\Big{(}m_{d}(parent(i))\geq\left\|\boldsymbol{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big{)},$
where any Gaussian that is sufficiently small at its current camera distance and whose parent is too large to be rendered, should be part of the cut set. This produces a proper cut under the assumption that child Gaussians always have a smaller minimal distance than their parents:
$\forall i:m_{d}(i)<m_{d}(parent(i)).$
This is generally valid when hierarchies are constructed after training, since parent Gaussians represent coarser approximations of their children. However, when the hierarchy is modified during training and densification, optimization can cause this assumption to break—leading to invalid cut sets and degenerate hierarchies that worsen over time.
@@
We present the hierarchical SPT data structure, which combines the benefits of both approaches. To construct it, we cut it using a BFS on the condition
$c_{HSPT}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$
with volume threshold size. The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
The volume of each root node in the lower hierarchy is now bounded by size, which also roughly bounds the extent of all Gaussians in the subtree. This provides an upper bound on the error introduced if the subtree is converted into an SPT. Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into a Sequential Point Tree to accelerate cut computation. The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view. Second, each selected SPT is cut according to the camera’s distance to its root node. Together, these yield the full set of Gaussians needed for rendering the current frame.
The construction and cutting process of an HSPT is illustrated in Figure 4. The SPTs used to render a MatrixCity-Scale frame are shown in Figure 11, and Figure 10 demonstrates the smooth LoD transitions enabled by the hierarchical structure.
Figure 4. A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
Rebuilding the hierarchical SPT for every training iteration would eliminate any performance benefit. Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently. In practice, we rebuild the HSPT only after each densification step.
@@
$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$
@@
The major benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees. To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum. Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in practice. It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
4.5. Caching on the GPU
Loading Gaussian data from RAM is a costly operation that can become a significant bottleneck during large-scale training. To mitigate this, we maintain a GPU-resident cache of Gaussians that are likely to be reused across consecutive training views. However, checking the cache for every individual Gaussian would introduce nontrivial overhead. Once again, SPTs offer an efficient alternative.
Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$th SPT. During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\boldsymbol{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is already cached. This check is performed using a simple distance ratio tolerance:
$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$
@@ 4.5. Caching on the GPU
To bound VRAM usage, the we use a least-recently-used (LRU) write-back policy. When a memory threshold is exceeded, entries are written back to RAM. Additionally, to prevent overfitting to persistent cache entries, the entire cache is flushed every 1 00010001\,0001 000 iterations. Figure 5 illustrates the caching process across two frames.
Figure 5. Caching strategy overview. Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits. After optimization, newly accessed SPTs are added to the GPU cache.
@@
In large-scale scenes, the GPU cache typically covers only a small fraction of the overall geometry, leading to sparse cache hits. To improve cache utilization, we prioritize spatial locality by selecting successive training views that are geographically close to the current one, thereby maximizing Gaussian reuse.
To this end, we precompute a $k$-nearest-neighbor graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$. The next training view $j$ is then sampled from the $k$-nearest neighbors of the current view $i$ according to the distribution:
$\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$
where $W$ is a constant, controlling the degree of exploration.
However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce bias. Thus, we inject a random view every 12 iterations, which we find sufficient to preserve generalization performance.
4.6. Memory Layout
Figure 6 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MatrixCity-Scale dataset. The output frame from this iteration is shown in Figure 12. The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states. In contrast, the hierarchy structure itself accounts for less than 10% of the total RAM footprint. On the GPU, the SPT metadata for all 60 million Gaussians occupies just 680 MB of VRAM and the upper hierarchy negligible 24 MB.
@@ 4.6. Memory Layout
The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as PyTorch overhead. Figure 15 provides a breakdown of memory usage for Gaussian properties and hierarchy data, and how these are distributed across CPU and GPU.
Figure 6. Peak memory consumption of CPU and GPU for a training iteration on BigCity-Scale with 60 million Gaussians.
5. Evaluation
Choice of dataset
Our method is designed to enable seamless training and rendering on large-scale scenes comprising thousands of views captured at vastly different scales. Unfortunately, most existing datasets of sufficient size contain either street-level or aerial views—but not both. Additionally, at this scale and density, current Structure-from-Motion (SfM) pipelines often fail to reliably recover camera poses and sparse point clouds.
To overcome these challenges, we use the synthetic MatrixCity dataset (Li et al., 2023), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry. For our main benchmark, MatrixCity-Scale (MC-Scale), we aggregate 11 466 street-view images, 3 208 aerial views, and 533 additional high-altitude views we generate manually. We evaluate reconstruction quality on a separate set of 1 694 test images. The diversity in image scale is visualized in Figure 8. To validate our method on real-world data, we also include results on the largest public dataset from Kerbl et al. (2024), the Campus scene.
Choice of Comparison
Selecting appropriate baselines is challenging. Standard 3DGS variants perform poorly at large scales, while most divide-and-conquer approaches—such as Hierarchical 3DGS, CityGS, and HorizonGS—assume access to 80 GB GPUs. Despite access to multiple clusters, we were limited to 40 GB A100s, highlighting the hardware demands of these methods.
To our knowledge, only Hierarchical 3DGS has trained successfully on the Campus dataset, and no prior work has demonstrated large-scale training on MatrixCity with combined aerial and street views. Due to resource constraints, we were unable to train CityGS on either dataset and adopt Hierarchical 3DGS as our main baseline.
For additional reference, we evaluate 3DGS-MCMC, modified to support large-scale training by limiting densification, streaming images from disk, and reducing the SH degree. Most chunk-based methods, including Kerbl et al. (2024), require feature correspondences, which are not provided in MatrixCity. To enable comparison, we reconstructed a sparse point cloud using COLMAP and the provided camera poses. This less stable point cloud (visualized in Figure 8) was used for initialization.
On the Campus scene, we report results for the official test split (Hierarchical single) and the full pretrained model. Consequently, we evaluate reconstruction quality on every 100th training view. We train a smaller model (38M) and a larger (80M) model. Rendering Hierarchical 3DGS required reducing SH degree to 1—even on 40 GB VRAM; ours also uses SH degree 1. Given the limitations of numerical metrics in these settings, we strongly encourage visual comparison via the supplementary videos, which in our opinion favour Hierarchical 3DGS. Except for the results of Hierarchical 3DGS, all training and rendering was done using an RTX 3090 GPU.
5.1. Results
Quantitative comparisons are shown in Table 1. A LoD of Gaussians successfully reconstructs the full MC-Scale scene, where other methods struggle due to the scene’s size and diverse camera viewpoints. Notably, Kerbl et al. (2024) reconstructs individual chunks accurately, but the merging process introduces ”floaters”—isolated Gaussians not associated with meaningful geometry—which obscure most test views (see Figure 13 and 14). Similarly, 3DGS-MCMC suffers from aggressive opacity and scale-based pruning, causing a large fraction of Gaussians to vanish entirely.
On the Campus dataset, Hierarchical 3DGS performs well, as its chunk-based approach aligns with the dataset’s single, street-level trajectory. Their single-chunk evaluation shows strong metrics. However, when evaluating their full model with reduced SH degree (to fit within 40 GB VRAM), we observe a clear drop in numerical quality. That said, the visual results remain competitive—see the supplemental video. In our case, both low- and high-resolution variants achieve similar quantitative scores despite visible differences, due to Hierarchical 3DGS evaluating on training views.
For context, Hierarchical 3DGS splits the MC-Scale scene into 94 chunks, each trained for 45 000 iterations, totaling over 4.23 million steps and nearly four full days of compute time on our machine. By contrast, our approach trains the full scene end-to-end in just 250 000 iterations, completing in approximately 15 hours.
Memory
A major limitation of chunk-based methods is the memory overhead during rendering. Although it also implements a form of GPU streaming, we were unable to run the viewer provided by Kerbl et al. (2024) on either dataset using the 40 GB VRAM A100, which is consistent with the limitations they report. In contrast, our method renders interactive flythroughs of MatrixCityScale and Campus (38M), see supplemental, while only requiring 8 GB of GPU memory (and 16 GB for Campus (80M)).
Table 1. Full Novel View Synthesis results. * use of the suboptimal COLMAP version for MC-Scale; adjusted SH degree to match ours and reduce memory to enable rendering, use of higher SH degree and only single chunk, reproduced from (Kerbl et al., 2024).
MC-Scale (15k images) Campus (22k images)

### v3

@@ 4 Method
To train models that exceed GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view according to the LoD hierarchy. To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees are replaced by Sequential Point Trees (SPTs), forming a hierarchical SPT (HSPT). To minimize costly transfers between RAM and VRAM, we track which SPTs currently reside in GPU memory and at which detail. Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM. Densification is performed on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes. The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations. Appendix A.2 includes more details on initialization and training. An overview of our training and densification process can be found in Figure 8 and pseudocodes in Appendix A.8.
4.1 Sequential Point Trees for Gaussian Splatting
Figure 1: An SPT and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red. Vertical lines in the SPT show the binary search result and horizontal lines the distance cut.
Sequential Point Trees (SPT) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians. They enforce a more constrained but more efficient cut condition than Hierarchical:
$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$ (4)
This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$. It requires storing only sorted pairs $\big(m_{d}(i),m_{d}(\text{parent}(i))\big)$, significantly reducing memory compared to full Gaussian hierarchies. To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order. A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered. Note that cuts are guaranteed to be proper, with nodes where $m_{d}(\text{parent}(i))>m_{d}(i)$ never being selected for the cut. Evidently, the level of detail of all Gaussians in the SPT is dictated by the camera’s distance to its root node. This can lead to Gaussians with $m_{d}(i)>\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view. To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function. By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$. In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail. SPTs are best suited for tightly grouped Gaussians observed from distances greater than their mutual spacing. Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes. Figure 1 visualizes both hierarchy types and their LoD cuts for a toy example.
4.2 Densification
Figure 2: Example of densifying and respawning leaf nodes.
@@ 4.2 Densification
Densifying an LoD representation presents a unique challenge, as the hierarchical structure must evolve continuously during training. Prior works circumvent this issue by constructing LoD hierarchies only after chunk-level training and densification are complete.
@@
We take inspiration from 3DGS-MCMC (3DGSasMCMC), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original. Notably, this mirrors how a parent node in a Gaussian hierarchy approximates its children. Therefore, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts.
@@
Instead of pruning, 3DGSasMCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian. We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified. See Figure 2 for an overview of the two hierarchy densification operations. Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required.
While 3DGSasMCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Hierarchical, which selects Gaussians for densification based on their maximal screen-space gradient. Avoiding destructive opacity resets, as advocated by 3DGS, is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal. Further details on our densification method can be found in Appendix A.7.
4.3 The hierarchical SPT datastructure
This section covers previous approaches to LoD selection and why a new datastructure—the hierarchical SPT—is necessary for robust and efficient training.
@@ 4.3 The hierarchical SPT datastructure
To enable GPU-accelerated cuts, Hierarchical evaluate the cut condition in parallel for each Gaussian:
$\Big(m_{d}(i)<\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(parent(i))\geq\left\|\bm{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big),$ (5)
where any Gaussian that is sufficiently small at its current camera distance and whose parent is too large to be rendered, should be part of the cut set. This produces a proper cut under the assumption that child Gaussians always have a smaller minimal distance than their parents (i.e. the heap condition is fulfilled): $\forall i:m_{d}(i)<m_{d}(parent(i)).$ This is generally valid when hierarchies are constructed after training, since parent Gaussians represent coarser approximations of their children. However, when the hierarchy is modified during training and densification, optimization can break the heap condition—leading to invalid cut sets and degenerate hierarchies that worsen over time.
@@
Figure 3: A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
We present the hierarchical SPT data structure, which combines the benefits of both approaches. To construct it from a Gaussian hierarchy, we cut it using a BFS on the condition $c_{\text{HSPT}}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$ with volume threshold size. The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
The volume of each root node in the lower hierarchy is now bounded by size, which also roughly bounds the extent of all Gaussians in the subtree. This provides an upper bound on the error introduced if the subtree is converted into an SPT. Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into a Sequential Point Tree to accelerate cut computation. The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view. Second, each selected SPT is cut according to the camera’s distance to its root node. Together, these yield the full set of Gaussians needed for rendering the current frame. The construction and cutting process of an HSPT is illustrated in Figure 3. Figure 10 and Figure 15 show example frames with respectively SPTs highlighted and different levels of detail.
Rebuilding the HSPT for every training iteration would eliminate any performance benefit. Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently. In practice, we rebuild the HSPT only after each densification step.
@@
$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$ (6)
@@
The major benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees. To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum. Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in our experiments. It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
4.4 Caching on the GPU
Figure 4: Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits. After optimization, newly accessed SPTs are added to the GPU cache.
Loading Gaussian data from RAM is a costly operation that can become a significant bottleneck during large-scale training. To mitigate this, we maintain a GPU-resident cache of Gaussians that are likely to be reused across consecutive training views. However, checking the cache for every individual Gaussian would introduce non-trivial overhead. Once again, SPTs offer an efficient alternative.
Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$-th SPT. During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\bm{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$ (7)
@@ 4.4 Caching on the GPU
To bound VRAM usage, we use a least-recently-used (LRU) write-back policy. When a memory threshold is exceeded, entries are written back to RAM. Additionally, to prevent overfitting to persistent cache entries, the entire cache is flushed every 1 0001\,000 iterations. Figure 4 illustrates the caching process across two frames.
@@
In large-scale scenes, the GPU cache typically covers only a small fraction of the overall geometry, leading to sparse cache hits. To improve cache utilization, we prioritize spatial locality by selecting successive training views close to the current one, maximizing Gaussian reuse.
To this end, we precompute a $k$-nearest-neighbour graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$. The next training view $j$ is then sampled from the $k$-nearest neighbours of the current view $i$ according to the distribution: $\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$ where $W$ is a normalization constant that also controls the degree of exploration.
However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce training bias. To counteract this, we inject a randomly selected view every 128 iterations, which we find sufficient to preserve generalization performance.
4.5 Memory Layout
Figure 5: Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
Figure 5 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MC-smaller-city+ dataset. The output frame from this iteration is shown in Figure 13. The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states. In contrast, the hierarchy structure itself accounts for less than 10% of the total RAM footprint. On the GPU, the SPT metadata for all 60 million Gaussians occupies just 680 MB of VRAM and the upper hierarchy negligible 24 MB.
@@ 4.5 Memory Layout
The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as general PyTorch overhead. Figure A.1 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
5 Evaluation
Our method is designed to enable seamless training and rendering on ultra-large-scale scenes comprising tens of thousands of views captured at vastly different scales. Unfortunately, most datasets of sufficient size contain either street-level or aerial views—but not both. To overcome these challenges, we turn to the synthetic MatrixCity dataset (MatrixCityDataset), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry. We also include results on the real-world street-level datasets from H-3DGS. As baseline, we compare to recent large-scale 3DGS methods CityGaussian (CityGaussian), H-3DGS (Hierarchical), HorizonGS (HorizonGS) and OctreeGS (OctreeGS). We note that HorizonGS is not evaluated on the H-3DGS dataset, since it requires aerial views for training. Further results and additional details on the datasets and baselines and are included in A.4. Evaluations were run on a single H200 GPU with 141 GB VRAM, while all of our models were trained on a single RTX 3090 graphics card with 24 GB VRAM. The contribution of our components is ablated in Appendix A.3.
5.1 Results
Qualitative comparisons are shown in Figure 6 and quantitative comparisons in Tables 1 and 2. The outputs of H-3DGS and HorizonGS show significant floating and ghosting artifacts. While H-3DGS performs well on individual chunks of MC-smaller-city+, oversized floaters survive the merging procedure and obscure the majority of test images (cf. Figure 11, 13 and 17). The merging procedure of CityGaussian (CityGaussian), which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, leading to mostly artifact-free views, but significantly blurry results. A LoD of Gaussians reconstructs novel views with a high degree of detail while avoiding chunk-based artifacts (cf. Figure 16), resulting in a considerable lead in quality metrics. Seamless training enables faster convergence during training, reducing the required number of training iterations considerably over divide-and-conquer based methods.
Table 1: MatrixCity novel view synthesis results. Results with * use the MC-smaller-city+ dataset due to memory constraints. Results with require suboptimal COLMAP initialization instead of the dataset’s provided point cloud. VRAM usage is measured while rendering the test images. Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during rendering or training.
MC-small-city+ (42.2k images) / *MC-smaller-city+ (15.1k images)